1. A STEM student is trying to estimate the average velocity of the players in a basketball team by studying the previous game of the players. One particular player can produce an initial velocity of 5 m/s . If the player throws a basketball from a height of 2 meters at an angle of 45 degrees with respect to the horizontal axis and the given average velocity of 5 m/s , calculate: a. The maximum height reached by the ball; b. The maximum horizontal range to reach the same level as it was thrown; and c. The time for the ball to reach its maximum height
1
Expert's answer
2020-11-03T10:31:25-0500

a) Let's write the equations of motion of the ball in horizontal and vertical directions:


"x=v_0tcos\theta, (1)""y=y_0+v_0tsin\theta-\dfrac{1}{2}gt^2, (2)"

here, "x" is the horizontal displacement of the ball (or the range of the ball), "v_0=5.0\ \dfrac{m}{s}" is the initial velocity of the ball, "t" is the total flight time of the ball, "\theta=45^{\circ}" is the launch angle, "y" is the vertical displacement of the ball (or the height), "y_0=2.0\ m" is the initial height from which the ball was thrown and "g=9.8\ \dfrac{m}{s^2}" is the acceleration due to gravity.

Let's first find the time that the ball takes to reach the maximum height from the kinematic equation:


"v_y=v_0sin\theta-gt_{rise},""0=v_0sin\theta-gt_{rise},""t_{rise}=\dfrac{v_0sin\theta}{g}."

Then, we can substitute "t_{rise}" into the second equation and find the maximum height:


"y_{max}=y_0+v_0sin\theta\cdot\dfrac{v_0sin\theta}{g}-\dfrac{1}{2}g(\dfrac{v_0sin\theta}{g})^2,""y_{max}=y_0+\dfrac{v_0^2sin^2\theta}{2g},""y_{max}=2.0\ m+\dfrac{(5.0\ \dfrac{m}{s})^2\cdot sin^245^{\circ}}{2\cdot 9.8\ \dfrac{m}{s^2}}=2.64\ m."

b) Let's first find the time the ball takes to reach the same level as it was thrown:


"t=2t_{rise}=\dfrac{2v_0sin\theta}{g}."

Then, we can substitute "t" into the first equation and find the maximum horizontal range to reach the same level as the ball was thrown:


"x=v_0cos\theta\cdot \dfrac{2v_0sin\theta}{g}=\dfrac{v_0^2sin2\theta}{g},""x=\dfrac{(5.0\ \dfrac{m}{s})^2\cdot sin2\cdot45^{\circ}}{9.8\ \dfrac{m}{s^2}}=2.55\ m."

c) Finally, we can find the time that the ball takes to reach the maximum height from the formula:


"t_{rise}=\dfrac{v_0sin\theta}{g}=\dfrac{5.0\ \dfrac{m}{s}\cdot sin45^{\circ}}{9.8\ \dfrac{m}{s^2}}=0.36\ s."

Answer:

a) "y_{max}=2.64\ m."

b) "x=2.55\ m."

c) "t_{rise}=0.36\ s."